fifty android tablet will be Delivered to a certain school for testing but five of them are defective.The principal will get two of these tablets for testing.Let X represent the defective tablet,Y for non defective tablet and Z represents the random variable for which a non defective tablet will occur.find the variables of Z

1
Expert's answer
2022-02-27T11:59:19-0500

"P(Z=0)=\frac{C_{45}^2C_5^0}{C_{50}^2}=\frac{198}{245}."

"P(Z=1)=\frac{C_{45}^1C_5^1}{C_{45}^2}=\frac{45}{245}."

"P(Z=2)=\frac{C_{45}^0C_5^2}{C_{50}^2}=\frac{2}{245}."